# Augmentation categories in higher dimensions

### v3

@@ 2.3. $A_{\infty}$-categories and augmented DGAs
In other words, a semi-free DGA is a DGA $\mathcal{A}$ equipped with a set $\mathcal{S}=\mathcal{R}\sqcup\mathcal{H}$ of generators, such that $\mathcal{A}$ is the result of taking the free non-commutative unital $\mathbb{Z}/2$-algebra generated by the elements of $\mathcal{S}$, and quotienting by relations on $\mathcal{H}\cup\{1\}$. In addition, for any $h\in\mathcal{H}$, we have $|h|=0$, and $\partial h=0$.
@@ 2.4. Link grading
Let $\mathcal{A}$ be a semi-free DGA with generating set $\mathcal{S}=\mathcal{R}\sqcup\mathcal{H}$. Let $P$ be a finite set. A $P$-link grading on $\mathcal{A}$ is a choice of a pair of maps
@@ 3. Construction and invariance of the augmentation category
In [NRSSZ], augmentation categories are constructed by using a consistent sequence of perturbations. We use a different approach, by choosing a set of perturbations with no consistency in mind, and then localise (roughly following the construction of wrapped Fukaya categories due to Abouzaid-Seidel, explained in [GPS1]*chapter 3).
@@ 3.1. An auxiliary category
Let $\{\mathcal{A}^{P}\}_{P\subset\mathbb{N}}$ be a set of semi-free DGAs $\mathcal{A}^{P}$ with generating sets $\mathcal{S}^{P}=\mathcal{R}^{P}\sqcup\mathcal{H}^{P}$, indexed by all finite subsets $P$ of $\mathbb{N}$. Assume each $\mathcal{A}^{P}$ is equipped with a $P$-link grading. We say that $\{\mathcal{A}^{P}\}_{P\subset\mathbb{N}}$ is a system of semi-free DGAs if:
@@
Let $(M,d\lambda)$ be a $2n$-dimensional exact symplectic manifold with finite geometry at infinity. Let $V=M\times\mathbb{R}$. Equip $V$ with the standard contact form $\alpha=dz-\lambda$, where $z$ is the $\mathbb{R}$ component. Let $\Lambda\subset V$ be a compact Legendrian submanifold with vanishing Maslov class. Let $\Lambda\subset U_{\Lambda}\subset V$ be a small contact neighborhood of $\Lambda$. We require that the contact neighborhood $U_{\Lambda}$ is small enough so that Theorem 2.9 applies. Pick a sequence of increasing small functions $f_{1}<f_{2}<\cdots$, and let $\Lambda^{i}:=\Gamma_{df_{i},\Lambda}\subset U_{\Lambda}\subset V$. Also pick an almost complex structure $J$ (guaranteed by Theorem 2.9) such that counting rigid holomorphic disks is equivalent to counting rigid quantum flow trees. We will refer to such an almost complex structure as compatible.
On each $\Lambda^{i}$, choose base points so that they are sent to the same point on $\Lambda$ via the base projection. Reeb chords of $\coprod_{i\in\mathbb{N}}\Lambda^{i}$ are partitioned into long Reeb chords (which are not contained in $U_{\Lambda}$) and short Reeb chords (which are contained in $U_{\Lambda}$). Long Reeb chords (and their endpoints) from $\Lambda^{i}$ to $\Lambda^{j}$ correspond to Reeb chords on $\Lambda$ (and their endpoints), and the base projection of the endpoints of a long Reeb chord from $\Lambda^{i}$ to $\Lambda^{j}$ onto $\Lambda$ are close to the corresponding endpoints of the Reeb chord on $\Lambda$. Choose endpoint paths so that for an endpoint of a Reeb chord on $\Lambda$, all the endpoint paths of the corresponding long Reeb chord endpoints are close to each other after base projection. Short Reeb chords from $\Lambda^{i}$ to $\Lambda^{j}$ (for $i,j$) correspond to critical points of $f_{j}-f_{i}$. Choose endpoint paths of a short Reeb chord from $\Lambda^{i}$ to $\Lambda^{j}$ so that the path for the beginning point and the path for the ending point project to the same path on $\Lambda$.
For generically chosen functions $f_{i}$ and compatible almost complex structure $J$, we have that for any finite set $P\subset\mathbb{N}$, the Legendrian $\Lambda^{P}:=\coprod_{i\in S}\Lambda^{i}$ is quantum flow tree-generic, and so gives rise to a CE-DGA which we denote by $\mathcal{A}^{P}$. By remark 2.16, $\mathcal{A}^{P}$ is equipped with a $P$-link grading. This gives rise to a system of semi-free DGAs $\{\mathcal{A}^{P}\}_{P\subset\mathbb{N}}$.
@@
Without vanishing Maslov class, the construction still goes through, just with $\mathbb{Z}/2$-grading on all the hom spaces.
@@
Without loss of generality, assume that $i_{1}<...<i_{k+1}$. In other cases, either the map is determined by strict unitality, or has trivial domain. Write $P=\{i_{1},...,i_{k+1}\}$. Consider the diagonal augmentation $\epsilon=(\epsilon_{1},...,\epsilon_{k+1})$ of $\mathcal{B}^{P}$, and let $f^{P}_{\epsilon}:=\phi_{\epsilon}\circ f^{P}\circ\phi^{-1}_{\epsilon\circ f^{P}}$. Here, we used that $f^{P}$ passes to a well-defined map $(\mathcal{A}^{P})^{\epsilon\circ f^{P}}\to(\mathcal{B}^{P})^{\epsilon}$, which can be shown by a direct computation. Another direct computation shows that $f^{P}_{\epsilon}((\mathcal{A}^{P})^{\epsilon\circ f^{P}}_{+})\subset(\mathcal{B}^{P})^{\epsilon}_{+}$, i.e., no constant terms appear in the image of generators. We then define $F_{k}$, up to the usual grading shift, by dualizing the component of $f^{S}_{\epsilon}$ that maps from
@@
Combining Definition 3.6 and Lemma 3.7, we obtain:
Corollary 3.8.
Let $\Lambda$ to $\Lambda^{\prime}$ be Legendrian isotopic. Let $f_{i}:\Lambda\to\mathbb{R}$, $f^{\prime}_{i}:\Lambda^{\prime}\to\mathbb{R}$. Let $J,J^{\prime}$ be compatible almost complex structures for the neighborhoods $U_{\Lambda},U_{\Lambda^{\prime}}$, then Definition 3.6 gives an $A_{\infty}$-functor $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$.
@@
The functor $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$ in Corollary 3.8 is an $A_{\infty}$-equivalence.
@@ 3.2. The augmentation category
The pre-augmentation category does not have enough morphisms: all objects corresponding to a single augmentation morally should be in the same isomorphism class. However, in the pre-augmentation category, there is no morphism from $(2,\epsilon)$ to $(1,\epsilon)$. To fix this, and to obtain the actual augmentation category, we localise. Remember that we required our perturbations to be small, so that all the $\mathcal{A}^{\{i\}}$ are canonically isomorphic, so their augmentations are in canonical bijection. For each choice of $i$, the Reeb chords of $\Lambda^{i}\cup\Lambda^{i+1}$ contain the set of short Reeb chords, corresponding to the critical points of a Morse function $f_{i+1}-f_{i}:\Lambda\to\mathbb{R}$. Define
@@
To prove that $w_{\epsilon,i}^{\vee}$ is a cocycle, we use quantum flow trees. By Lemma 2.11, the rigid holomorphic disks that have a negative puncture at some $p_{l}$ are rigid complete flow lines of $f_{i+1}-f_{i}$ that start from some index 1 critical point and end at $p_{l}$, and multiscale flow trees consisting of a constant disk with a partial flow line that flows to $p_{l}$.
The contribution from the former is 0, since the sum of all minimums of a Morse function is a cocycle in the Morse cochain complex.
The contribution from the latter rigid quantum flow trees cancel out in pairs: for each Reeb chord $b$ of $\Lambda$, there are four corresponding long Reeb chords in $\Lambda^{\{i,i+1\}}$, denoted $b_{i,i},b_{i,i+1},b_{i+1,i},b_{i+1,i+1}$ (the first subscript is the number of the beginning sheet, and the second subscript is the number of the ending sheet). Since we are only interested in the differential in $C_{i,i+1}^{\vee}$, we only look at rigid holomorphic disks with a positive puncture at $b_{i,i+1}$. Given a quantum flow tree consisting of a constant disk at $b$ and a Morse flow line that flows to some minimum, there are two lifts to a disk. One consists of a negative puncture at a minimum and a negative puncture at $b_{i,i}$; the other one consists of a negative puncture at a minimum and a negative puncture at $b_{i+1,i+1}$. Both these rigid quantum flow trees contribute the summand $\epsilon(b)b_{i,i+1}^{\vee}$ to $m_{1}(\epsilon)(w_{\epsilon,i}^{\vee})$, so they cancel out and contribute 0.
@@
The cocycle $w_{\epsilon,i}^{\vee}$ has grading 0 since we obtain the grading by picking the same basepoint on both sheets, and pick the same path to the beginning and end point of a short Reeb chord, which completes to a trivial loop in the Lagrangian Grassmanian, since it is constructed as a path concatenated with the inverse of the path. ∎
@@
Let $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$ be the functor defined in Corollary 3.8. Then,
@@
Let $f,g:\Lambda\to\mathbb{R}$ be small positive functions. Choose
@@
Perturb $f,g$ so that $\coprod_{i=1,2,3}\Lambda^{i}$ and $\coprod_{i=1,2,3}\Lambda^{\prime i}$ are both quantum flow tree-generic. If $\epsilon$ is small enough, then throughout the straight line homotopy $f^{t}_{i}$ from $f_{i}$ to $f^{\prime}_{i}$ (where only $f_{2}$ is changing), we have that the functions
@@
are up to scaling, all small perturbations of the function $f$. We have that throughout the homotopy for $1\leq i,j\leq 3$, we have that
@@
We have that
$H^{*}m_{2}:H^{*}\mathcal{O}_{\Lambda,f_{i},J}((\epsilon,1),(\epsilon,2))\otimes H^{*}\mathcal{O}_{\Lambda,f_{i},J}((\epsilon,2),(\epsilon,3))\to H^{*}\mathcal{O}_{\Lambda,f_{i},J}((\epsilon,1),(\epsilon,3))$
gives rise to a multiplication map
@@
A priori, if we replace $f_{i}$ by $f^{\prime}_{i}$, we might obtain different multiplication maps. However, as we have shown above, the straight line homotopy between the 2 choices induces canonical isomorphisms of groups between the 2 choices. We then use the fact that $H^{*}m_{2}$ is natural with respect to different choices of $f_{i}$ to conclude that the 2 choices give the same multiplication map $M$. Note that the multiplication map only depends on $f_{2}-f_{1}$, and it does not depend on $f_{i}$ for any other $i$.
@@
$w_{\epsilon,1}^{\vee}]\in H^{0}\mathcal{O}((\epsilon,1),(\epsilon,2))$ is the cohomological unit for $M$.
@@
Now we compute the product $M(\cdot,[w_{\epsilon,1}^{\vee}])$ using quantum flow trees. For quantum flow trees here, we use Corollary 2.12, and we degenerate $\Lambda^{\{1,2,3\}}$ onto $\Lambda^{\{2,3\}}$, since $\Lambda^{1}$ and $\Lambda^{2}$ are close together.
@@
For a Reeb chord $b_{23}$, when computing $m_{2}(b_{23}^{\vee},w_{\epsilon,1}^{\vee})$, we count rigid quantum flow trees with a negative puncture at some $(p_{l})_{12}$, and a negative puncture at $b_{23}$. There is precisely one such rigid quantum flow tree, which is a constant disk at $b$, with a small flow line flowing down to a particular $p_{l}$. After lifting the quantum flow tree, we get a disk with positive puncture at $b_{13}$, and negative punctures at $(p_{l})_{23}$ and $b_{12}$. So we have that $m_{2}(b_{23}^{\vee},w_{\epsilon,12}^{\vee})=b_{13}$. So, $(w_{\epsilon,1}^{\vee})$ is a right unit.
@@
The multiplication $M$ is a natural with respect to Legendrian isotopies. In other words,

### v4

@@ 2.3. $A_{\infty}$-categories and augmented DGAs
In other words: let $R$ be a non-commutative coefficient ring, with a generating set $\mathcal{H}$, and some relations. A semi-free DGA is a DGA $\mathcal{A}$ equipped with a set $\mathcal{S}=\mathcal{R}\cup\mathcal{H}$ of generators, such that $\mathcal{A}$ is the result of taking the free non-commutative unital $\mathbb{Z}/2$-algebra generated by the elements of $\mathcal{S}$, and quotienting by relations on $\mathcal{H}$ that are inherited from the relations that define $R$. In addition, for any $h\in\mathcal{H}$, we have $|h|=0$, and $\partial h=0$.
@@ 2.4. Link grading
Let $\mathcal{A}$ be a semi-free DGA with generating set $\mathcal{S}=\mathcal{R}\cup\mathcal{H}$. Let $P$ be a finite set. A $P$-link grading on $\mathcal{A}$ is a choice of a pair of maps
@@ 3. Construction and invariance of the augmentation category
In [NRSSZ], augmentation categories are constructed by using a consistent sequence of perturbations. Since finding a consistent sequence of perturbations in higher dimensions is difficult, we use a different approach, by choosing a set of arbitrary generic perturbations with no consistency in mind, and then localise (roughly following the construction of wrapped Fukaya categories due to Abouzaid-Seidel, explained in [GPS1]*chapter 3).
An advantage of using the localisation approach is that we get a more general invariance statement than using specific consistent sequences: suppose we find one family of consistent sequences of perturbations, and prove the invariance of the augmentation category obtained. If we find another family of consistent sequences, we do not automatically know if the new family of consistent sequences give rise to an equivalent augmentation category. Using the localisation approach, we prove that any arbitrary generic perturbations give rise to the same augmentation category, up to equivalence.
@@ 3.1. An auxiliary category
First, we build an auxiliary category (Definition 3.3), and prove its invariance (Proposition 3.9). The auxiliary category is constructed by considering countably many generic positive pushoffs $\Lambda^{i}$ of $\Lambda$. The objects of the auxiliary category are augmentations of some $\Lambda^{i}$, and the morphisms only “go up” in the index $i$. This has the advantage that we can perturb the infinitely many copies of $\Lambda$ generically, and obtain transversality of moduli spaces of holomorphic disks for any finite subset of them for free, by citing [EES2]. We start with some algebraic constructions.
@@
Let $\{\mathcal{A}^{P}\}_{P\subset\mathbb{N}}$ be a set of semi-free DGAs $\mathcal{A}^{P}$ with generating sets $\mathcal{S}^{P}=\mathcal{R}^{P}\cup\mathcal{H}^{P}$, indexed by all finite subsets $P$ of $\mathbb{N}$. Assume each $\mathcal{A}^{P}$ is equipped with a $P$-link grading. We say that $\{\mathcal{A}^{P}\}_{P\subset\mathbb{N}}$ is a system of semi-free DGAs if:
@@
Let $(M,d\lambda)$ be a $2n$-dimensional exact symplectic manifold with finite geometry at infinity. Let $V=M\times\mathbb{R}$. Equip $V$ with the standard contact form $\alpha=dz-\lambda$, where $z$ is the $\mathbb{R}$ component. Let $\Lambda\subset V$ be a compact Legendrian submanifold with vanishing Maslov class. Let $\Lambda\subset U_{\Lambda}\subset V$ be a small contact neighborhood of $\Lambda$. We require that the contact neighborhood $U_{\Lambda}$ is small enough so that Theorem 2.9 applies. Pick a sequence of increasing small functions $f_{1}<f_{2}<\cdots$, and let $\Lambda^{i}:=\Gamma_{df_{i},\Lambda}\subset U_{\Lambda}\subset V$. Also pick an almost complex structure $J$ (guaranteed by Theorem 2.9) such that counting rigid holomorphic disks is equivalent to counting rigid quantum flow trees.
On $\Lambda$, we pick a set of basepoints $B=\{p_{1},\ldots,p_{r}\}$, which contains one basepoint for each component of $\Lambda$, and one of the basepoints $p_{1}$ is designated to be the special basepoint. As in the setting before Definition 2.1, we pick paths $\gamma_{1j}$ connecting each $\Pi(p_{j})$ to $\Pi(p_{1})$, and a symplectic trivialisation along these paths that identify the tangent spaces of their endpoints. On $\Lambda^{i}$, pick $p_{j}^{i}$ to be a basepoint that is identified with $p_{j}$ on $\Lambda$ via the base projection map $\Lambda^{i}\hookrightarrow U_{\Lambda}\to\Lambda$. On $\coprod_{i\in\mathbb{N}}\Lambda^{i}$, we designate $p_{1}^{1}$ to be the special basepoint. We choose the curves that connect the basepoints to the special basepoint in the following way: connect all $\Pi(p_{1}^{i})$ to $\Pi(p_{1}^{1})$ via paths that are $C^{0}$-small. For another basepoint of $p_{i}^{j}$, we connect $\Pi(p_{i}^{j})$ to $\Pi(p_{1}^{j})$ by a path whose corresponding path of Lagrangian subspaces is $C^{0}$-close to the path of Lagrangian subspaces corresponding to $\gamma_{1j}$, and concatenate with the path from $\Pi(p_{1}^{j})$ to $\Pi(p_{1}^{1})$ that we picked above.
Reeb chords of $\coprod_{i\in\mathbb{N}}\Lambda^{i}$ are partitioned into long Reeb chords (which are not contained in $U_{\Lambda}$) and short Reeb chords (which are contained in $U_{\Lambda}$). Long Reeb chords (and their endpoints) from $\Lambda^{i}$ to $\Lambda^{j}$ correspond to Reeb chords (and their endpoints) on $\Lambda$, and the base projection of the endpoints of a long Reeb chord from $\Lambda^{i}$ to $\Lambda^{j}$ onto $\Lambda$ are close to the corresponding endpoints of the Reeb chord on $\Lambda$. Choose endpoint paths so that for an endpoint of a Reeb chord on $\Lambda$, all the endpoint paths of the corresponding long Reeb chord endpoints are close to each other after base projection. Short Reeb chords from $\Lambda^{i}$ to $\Lambda^{j}$ (for $i,j$) correspond to critical points of $f_{j}-f_{i}$. Choose endpoint paths of a short Reeb chord from $\Lambda^{i}$ to $\Lambda^{j}$ so that the path for the beginning point and the path for the ending point project to the same path on $\Lambda$.
For generically chosen functions $f_{i}$ and compatible almost complex structure $J$, we have that for any finite set $P\subset\mathbb{N}$, the Legendrian $\Lambda^{P}:=\coprod_{i\in S}\Lambda^{i}$ is quantum flow tree-generic, and so gives rise to a Chekanov-Eliashberg DGA which we denote by $\mathcal{A}^{P}$. By remark 2.16, $\mathcal{A}^{P}$ is equipped with a $P$-link grading. This gives rise to a system of semi-free DGAs $\{\mathcal{A}^{P}\}_{P\subset\mathbb{N}}$.
@@
Without vanishing Maslov class, the construction still goes through, with $\mathbb{Z}/2m_{\Lambda}$-grading on all the hom spaces, where $m_{\Lambda}$ is the gcd of the minimal Maslov numbers of all connected components of $\Lambda$.
@@
Without loss of generality, assume that $i_{1}<...<i_{k+1}$. In other cases, either the map is determined by strict unitality, or has trivial domain. Write $P=\{i_{1},...,i_{k+1}\}$. Consider the diagonal augmentation $\epsilon=(\epsilon_{1},...,\epsilon_{k+1})$ of $\mathcal{B}^{P}$, and let $f^{P}_{\epsilon}:=\phi_{\epsilon}\circ f^{P}\circ\phi^{-1}_{\epsilon\circ f^{P}}$. Here, we used that $f^{P}$ passes to a well-defined map $(\mathcal{A}^{P})^{\epsilon\circ f^{P}}\to(\mathcal{B}^{P})^{\epsilon}$, which can be shown by a direct computation. Another direct computation shows that $f^{P}_{\epsilon}((\mathcal{A}^{P})^{\epsilon\circ f^{P}}_{+})\subset(\mathcal{B}^{P})^{\epsilon}_{+}$, i.e., no constant terms appear in the image of generators. We then define $F_{k}$, up to the usual degree $1$ grading shift ($C^{\vee}=C^{*}[-1]$, from Section 2.3), by dualizing the component of $f^{S}_{\epsilon}$ that maps from
@@
Combining Definition 3.6 and Lemma 3.7, we can define:
Definition 3.8.
Let $\Lambda$ to $\Lambda^{\prime}$ be Legendrian isotopic. Let $f_{i}:\Lambda\to\mathbb{R}$, $f^{\prime}_{i}:\Lambda^{\prime}\to\mathbb{R}$. Let $J,J^{\prime}$ be compatible almost complex structures for the neighborhoods $U_{\Lambda},U_{\Lambda^{\prime}}$. Define $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$ as follows: since Lemma 3.7 guarantees us a set of compatible DGA morphisms, we can use Definition 3.6 to define an $A_{\infty}$-functor $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$.
@@
The functor $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$ in Definition 3.8 is an $A_{\infty}$-equivalence.
@@ 3.2. The augmentation category
Remember that in the construction before Definition 3.3, we required our perturbations to be small, so all the $\mathcal{A}^{\{i\}}$ are canonically isomorphic, so their augmentations are in canonical bijection. The pre-augmentation category does not have enough morphisms: since we set out to construct a category whose objects are augmentations of $\Lambda$, all objects corresponding to a single augmentation morally should be in the same isomorphism class. However, in the pre-augmentation category, there is no morphism from $(2,\epsilon)$ to $(1,\epsilon)$. To fix this, and to obtain the actual augmentation category, we localise. For each choice of $i$, the Reeb chords of $\Lambda^{i}\cup\Lambda^{i+1}$ contain the set of short Reeb chords, corresponding to the critical points of a Morse function $f_{i+1}-f_{i}:\Lambda\to\mathbb{R}$. Define
@@
To prove that $w_{\epsilon,i}^{\vee}$ is a cocycle, we use quantum flow trees. For any index 1 critical point $q$ of $f_{i+1}-f_{i}$, let $\gamma_{q}$ denote the loop on $\Pi(\Lambda)$ formed by the capping path from the basepoint to $q$, the flow line from $q$ to an index 0 critical point $p$, and the inverse of the capping path from the basepoint to $p$. By Lemma 2.11, the rigid holomorphic disks that have a negative puncture at some $p_{l}$ are rigid complete flow lines of $f_{i+1}-f_{i}$ that start from some index 1 critical point and end at $p_{l}$, and multiscale flow trees consisting of a constant disk with a partial flow line that flows to $p_{l}$.
For rigid complete flow lines, note that each such flow line contributes $\epsilon(\gamma_{p})\epsilon(\gamma_{p}^{-1})=1$. The total contribution is 0, since the sum of all minimums of a Morse function is a cocycle in the Morse cochain complex.
The contribution from the latter rigid quantum flow trees cancel out in pairs: for each Reeb chord $b$ of $\Lambda$, there are four corresponding long Reeb chords in $\Lambda^{\{i,i+1\}}$, denoted $b_{i,i},b_{i,i+1},b_{i+1,i},b_{i+1,i+1}$ (the first subscript is the number of the beginning sheet, and the second subscript is the number of the ending sheet). Since we are only interested in the differential in $C_{i,i+1}^{\vee}$, we only look at rigid holomorphic disks with a positive puncture at $b_{i,i+1}$. Given a quantum flow tree consisting of a constant disk at $b$ and a Morse flow line that flows to some minimum, there are two lifts to a disk. One consists of a negative puncture at a minimum $p$ and a negative puncture at $b_{i,i}$; the other one consists of a negative puncture at a minimum and a negative puncture at $b_{i+1,i+1}$. Let $\gamma_{b}$ denote the loop on $\Pi(\Lambda)$ formed by the capping path from the basepoint to $p$, the inverse of the partial flow line from the positive endpoint of $b$ to $p$, and the inverse of the capping path from the basepoint to the positive endpoint of $b$. Also, let $\gamma_{b}^{\prime}$ denote the loop on $\Lambda$ formed by the capping path from the basepoint to the negative endpoint of $b$, the partial flow line from the negative endpoint of $b$ to $p$, and the inverse of the capping path from the basepoint to $p$. When computing $m_{1}(\epsilon)(w_{\epsilon,i}^{\vee})$, one of these rigid quantum flow trees contribute the summand $\epsilon(b)\epsilon(\gamma_{b})\epsilon(\gamma_{b}^{-1})b_{i,i+1}^{\vee}=\epsilon(b)b_{i,i+1}^{\vee}$, and the other rigid quantum flow tree contributes the summand $\epsilon(b)\epsilon(\gamma_{b}^{\prime})\epsilon({\gamma_{b}^{\prime}}^{-1})b_{i,i+1}^{\vee}=\epsilon(b)b_{i,i+1}^{\vee}$, so they cancel out and contribute 0.
@@
We also need to prove that the index 0 critical points $p_{l}$ have degree 0. The construction before Definition 3.3 specifies choices of paths connecting short Reeb chord endpoints to the basepoints and the paths that connect different basepoints. We compute the grading by assembling these paths into a loop in $M$, and look at the corresponding loop in the Lagrangian Grassmanian. The loop in $M$ we get is the concatenation of: a path $\gamma$ from $\Pi(p_{l})$ to $\Pi(b_{j}^{i})$ (where $b_{j}^{i}$ is the basepoint of the component of $p_{l}$ on $\Lambda^{i}$), a path from $\Pi(b_{j}^{i})$ to $\Pi(b_{1}^{i})$ that is $C^{0}$-close to $\gamma_{1j}$, a path from $\Pi(b_{1}^{i})$ to $\Pi(b_{1}^{1})$ that is $C^{0}$ small, the inverse of the path from $\Pi(b_{1}^{i+1})$ to $\Pi(b_{1}^{1})$ that is $C^{0}$ small, a path from $\Pi(b_{1}^{i+1})$ to $\Pi(b_{j}^{i+1})$ that is $C^{0}$-close to the inverse of $\gamma_{1j}$, and the inverse of the path from $\Pi(p_{l})$ to $\Pi(b_{j}^{i+1})$ (which is $C^{0}$-close to the inverse of the path from $\Pi(p_{l})$ to $\Pi(b_{j}^{i})$). Whenever two paths are $C^{0}$-close above, we have also arranged in the construction so that the corresponding paths of Lagrangian subspaces are also $C^{0}$-close. We see that the loop in $M$, and the corresponding loop in the Lagrangian Grassmanian obtained in the end are contractible loops, therefore the Maslov class is 0. The degree of $p_{l}$ is therefore 01=10-1=-1 in the Chekanov-Eliashberg DGA. Since there is a grading shift by 1 degree going from the DGA to $C^{\vee}$ (see Section 2.3), the degree of $p_{l}$ in $\mathcal{O}_{\Lambda,f_{i},J}((i,\epsilon),(i+1,\epsilon))$ is 0. ∎
@@
Let $F:\mathcal{O}_{\Lambda^{\prime},f^{\prime}_{i},J^{\prime}}\to\mathcal{O}_{\Lambda,f_{i},J}$ be the functor defined in Definition 3.8. Then,
@@
Let $f,h:\Lambda\to\mathbb{R}$ be small positive functions. Choose
@@
Perturb $f,h$ so that $\coprod_{i=1,2,3}\Lambda^{i}$ and $\coprod_{i=1,2,3}\Lambda^{\prime i}$ are both quantum flow tree-generic. If $\epsilon$ is small enough, then throughout the straight line homotopy $f^{t}_{i}$ from $f_{i}$ to $f^{\prime}_{i}$ (where only $f_{2}$ is changing), we have that the functions
@@
are up to scaling, all small perturbations of the function $f$. Throughout the homotopy, for $1\leq i,j\leq 3$, we have that
@@
Taking cohomology, $m_{2}$ gives a map
$H^{*}m_{2}:H^{*}\mathcal{O}_{\Lambda,f_{i},J}((\epsilon,1),(\epsilon,2))\otimes H^{*}\mathcal{O}_{\Lambda,f_{i},J}((\epsilon,2),(\epsilon,3))\to H^{*}\mathcal{O}_{\Lambda,f_{i},J}((\epsilon,1),(\epsilon,3)).$
Applying the stable tame isomorphism invariance of the Chekanov-Eliashberg DGA from [EES2] to the DGA of the 3-copy Legendrian $\Lambda^{1}\cup\Lambda^{2}\cup\Lambda^{3}$, we have that $m_{2}$ is natural with respect to Legendrian isotopies, up to quasi-isomorphism. So, $H^{*}m_{2}$ is natural with respect to Legendrian isotopies. This gives rise to a multiplication map
@@
A priori, if we replace $f_{i}$ by $f^{\prime}_{i}$, we might obtain different multiplication maps. However, as we have shown above, the straight line homotopy between the 2 choices induces canonical isomorphisms of groups between the 2 choices. We then use the fact that $H^{*}m_{2}$ is natural with respect to different choices of $f_{i}$ (which can be realised by Legendrian isotopies) to conclude that the 2 choices give the same multiplication map $M$. Note that the multiplication map only depends on $f_{2}-f_{1}$, and it does not depend on $f_{i}$ for any other $i$.
@@
$w_{\epsilon,1}^{\vee}]\in H^{0}\mathcal{O}((\epsilon,1),(\epsilon,2))$ is the unit for $M$.
@@
Now we compute the product $M(\cdot,[w_{\epsilon,1}^{\vee}])$ using quantum flow trees. For quantum flow trees here, we use Corollary 2.12, and we degenerate $\Lambda^{\{1,2,3\}}$ onto $\Lambda^{\{2,3\}}$ (see the discussion after Corollary 2.12), since $\Lambda^{1}$ and $\Lambda^{2}$ are close together.
@@
For a Reeb chord $b_{23}$, when computing $m_{2}(b_{23}^{\vee},w_{\epsilon,1}^{\vee})$, we count rigid quantum flow trees with a negative puncture at some $(p_{l})_{12}$, and a negative puncture at $b_{23}$. Since we are composing a minimum $p_{l}$ on the right of $b_{23}^{\vee}$, we only consider quantum flow trees with a partial flow line to the right of $b_{23}^{\vee}$. Since the positive endpoint of $b$ flows to a unique minimum point $p_{l}$, there is precisely one such rigid quantum flow tree, which is a constant disk at $b$, with a small flow line flowing down to a particular $p_{l}$. After lifting the quantum flow tree, we get a disk with positive puncture at $b_{13}$, and negative punctures at $(p_{l})_{23}$ and $b_{12}$. So we have that $m_{2}(b_{23}^{\vee},w_{\epsilon,12}^{\vee})=b_{13}$ (similarly to the computation of $m_{1}(w_{\epsilon,1}^{\vee})$ in the proof of Lemma 3.10, the contributions from the homology classes cancel out). So, $(w_{\epsilon,1}^{\vee})$ is a right unit.
@@
Applying the stable tame isomorphism invariance of the Chekanov-Eliashberg DGA from [EES2] to the DGA of the 3-copy Legendrian $\Lambda^{1}\cup\Lambda^{2}\cup\Lambda^{3}$, we have that $m_{2}$ is natural with respect to Legendrian isotopies, up to quasi-isomorphism. This implies that the multiplication $M$ is natural with respect to Legendrian isotopies. In other words,